# Regularized Estimation of Loading Matrix in Factor Models for High-Dimensional Time Series

### v2

@@ 5 Simulation Study
[…]
Datasets are simulated from model (2). The nonzero elements in ${\mathbf{A}}^{s}$ are simulated from a truncated standard normal distribution with absolute values bounded above by 0.1. We set $r=3$, and $\mathbf{f}_{t}^{s}$ is generated from three independent AR(1) processes with an AR coefficient of $0.9$ and an innovation variance of 1. The diagonal elements of $\boldsymbol{\Sigma}_{e,t}$ are all 1, and its off-diagonal elements are 0.5. The number of factors is assumed to be known. For each setting, we generate 300 samples and compare the estimation results for the loading space and the number of nonzero loadings.
@@ 5.1 Study on $\delta$, $p$ and $n$
In this example, we consider different combinations of $p$ and $n$, where $p=20,50,100,200,500$ and $n=100,200,500$. In ${\mathbf{A}}^{s}$, the first $0.4p$ elements in the first column, the middle $0.4p$ elements in the second column, and the last $0.4p$ elements in the third column are nonzero, while all other elements are zero. Note that the second column of ${\mathbf{A}}^{s}$ contains nonzero elements that overlap with those in both the first and third columns.
Table 1 shows the average estimation error of the loading space for different methods with $\delta=0$ and $\delta=0.25$. Our proposed estimator performs well in estimating the loading space and outperforms others, especially when the sample size is small. It is interesting that when the threshold for the varimax is appropriately selected, varimax performs better than the method of Lam et al., (2011), which does not account for the sparsity of the loading matrix. However, the optimal threshold value depends on $p$, $n$, and $\delta$, so selecting the threshold value is particularly challenging when analyzing real data.
@@
$\delta$ $n$ $p$ eigen varimax1 varimax2 SO sparse
0 100 20 0.115(0.056) 0.115(0.056) 0.107(0.057) 0.187(0.131) 0.095(0.057)
100 50 0.110(0.038) 0.109(0.038) 0.103(0.033) 0.159(0.101) 0.088(0.036)
100 100 0.106(0.031) 0.103(0.032) 0.120(0.019) 0.134(0.068) 0.084(0.028)
100 200 0.103(0.028) 0.093(0.030) 0.174(0.012) 0.126(0.070) 0.082(0.024)
100 500 0.104(0.028) 0.085(0.028) 0.310(0.011) 0.125(0.066) 0.084(0.024)
200 20 0.066(0.022) 0.066(0.023) 0.066(0.021) 0.143(0.109) 0.052(0.023)
200 50 0.062(0.016) 0.060(0.017) 0.076(0.013) 0.091(0.053) 0.047(0.014)
200 100 0.061(0.015) 0.056(0.016) 0.105(0.010) 0.078(0.031) 0.046(0.013)
200 200 0.061(0.015) 0.050(0.015) 0.164(0.010) 0.076(0.025) 0.046(0.012)
200 500 0.061(0.014) 0.049(0.010) 0.303(0.009) 0.072(0.020) 0.046(0.012)
500 20 0.035(0.010) 0.035(0.010) 0.047(0.013) 0.098(0.067) 0.026(0.010)
500 50 0.035(0.008) 0.033(0.008) 0.066(0.012) 0.063(0.032) 0.024(0.007)
500 100 0.035(0.008) 0.032(0.008) 0.099(0.010) 0.051(0.020) 0.024(0.006)
500 200 0.035(0.007) 0.029(0.006) 0.159(0.009) 0.044(0.013) 0.024(0.006)
500 500 0.035(0.007) 0.035(0.004) 0.301(0.009) 0.042(0.012) 0.024(0.006)
0.25 100 20 0.152(0.074) 0.152(0.074) 0.143(0.076) 0.313(0.156) 0.130(0.077)
100 50 0.164(0.063) 0.163(0.063) 0.147(0.065) 0.386(0.161) 0.139(0.065)
100 100 0.170(0.053) 0.168(0.054) 0.159(0.046) 0.425(0.148) 0.145(0.053)
100 200 0.178(0.051) 0.173(0.053) 0.204(0.033) 0.477(0.130) 0.154(0.049)
100 500 0.203(0.058) 0.191(0.062) 0.334(0.023) 0.551(0.070) 0.181(0.055)
200 20 0.087(0.030) 0.087(0.030) 0.082(0.030) 0.264(0.150) 0.072(0.031)
200 50 0.091(0.023) 0.089(0.024) 0.089(0.018) 0.266(0.151) 0.073(0.022)
200 100 0.097(0.024) 0.093(0.025) 0.117(0.014) 0.331(0.169) 0.078(0.021)
200 200 0.106(0.025) 0.096(0.027) 0.174(0.012) 0.429(0.159) 0.087(0.021)
200 500 0.118(0.028) 0.098(0.029) 0.312(0.011) 0.520(0.110) 0.100(0.023)
500 20 0.046(0.013) 0.046(0.013) 0.053(0.014) 0.196(0.130) 0.036(0.014)
500 50 0.051(0.012) 0.049(0.012) 0.071(0.012) 0.199(0.133) 0.038(0.011)
500 100 0.056(0.013) 0.050(0.013) 0.104(0.011) 0.226(0.141) 0.042(0.011)
500 200 0.060(0.013) 0.049(0.013) 0.163(0.009) 0.298(0.168) 0.046(0.011)
500 500 0.067(0.014) 0.053(0.010) 0.304(0.009) 0.469(0.142) 0.052(0.011)
Table 2 shows the average value of $|m-\widehat{m}|$, with the standard deviation given in parentheses, where $\widehat{m}$ is the number of nonzero elements in the estimated loading matrix. Overall, our proposed method effectively recovers the sparsity of the loading matrix. The performance of “sparse” is much better than other methods. Although varimax rotation may be able to estimate the loading space accurately when the threshold is appropriately selected, it cannot recover the sparsity as effectively as our proposed method.
@@
Table 2: Mean and standard deviation (in parentheses) of $|m-\widehat{m}|$ for the example in Section 5.1
@@
$n$ $p$ varimax1 varimax2 SO sparse varimax1 varimax2 SO sparse
100 20 26.1(3.5) 4.4(3.2) 24.1(6.2) 2.6(3.1) 27.7(3.2) 6.1(3.9) 23.5(6.0) 3.4(3.8)
100 50 49.6(9.1) 5.3(2.9) 58.6(15.3) 5.0(6.9) 58.8(8.1) 5.2(4.9) 58.3(13.7) 7.0(8.3)
100 100 65.8(20.1) 22.2(4.1) 110.5(35.7) 7.8(12.3) 96.0(18.9) 20.5(5.3) 113.8(26.0) 12.9(19.3)
100 200 63.3(34.9) 72.8(6.1) 212.3(70.7) 10.7(14.8) 140.2(39.4) 72.4(7.4) 228.1(55.4) 21.5(28.6)
100 500 31.0(40.7) 300.1(9.3) 507.5(186.1) 19.3(29.9) 195.3(103.0) 302.6(9.3) 575.7(106.9) 64.2(70.9)
200 20 22.2(4.5) 2.6(2.5) 25.4(6.0) 2.3(3.5) 24.2(4.1) 3.2(2.8) 25.5(5.4) 3.0(3.9)
200 50 33.9(10.6) 6.2(2.6) 62.3(16.0) 2.2(3.3) 44.5(9.9) 5.8(2.6) 59.4(14.4) 4.0(6.0)
200 100 34.9(17.3) 22.6(3.8) 124.2(35.1) 3.2(5.4) 62.1(19.7) 22.6(3.9) 120.4(28.7) 5.9(10.0)
200 200 18.0(16.8) 72.4(6.4) 252.0(70.8) 5.6(9.2) 68.2(34.9) 72.6(6.3) 251.3(55.3) 11.4(20.2)
200 500 23.9(6.0) 298.6(9.7) 611.5(173.5) 9.6(13.8) 41.5(49.8) 299.6(9.7) 641.5(97.2) 26.8(31.0)
500 20 18.2(5.3) 2.0(2.0) 28.1(5.2) 1.4(2.9) 20.2(4.9) 2.2(2.1) 28.0(4.2) 2.1(3.7)
500 50 22.0(9.4) 6.1(2.5) 69.0(15.3) 1.2(3.6) 29.5(9.5) 6.0(2.6) 67.2(13.0) 2.5(5.3)
500 100 17.1(12.4) 23.0(4.0) 136.8(33.8) 1.4(3.8) 29.5(14.8) 23.0(4.1) 131.8(27.1) 3.3(7.9)
500 200 6.1(8.3) 72.1(6.3) 266.3(74.4) 2.1(5.2) 16.6(15.7) 72.3(6.5) 259.8(62.6) 5.1(13.4)
500 500 24.6(4.5) 299.0(9.4) 679.6(182.9) 5.2(13.0) 23.6(6.1) 299.5(9.5) 681.3(118.4) 7.9(18.4)
@@ 5.2 Study on $m$
In this section, we evaluate the impact of $m$ (the total number of nonzero loadings in ${\mathbf{A}}^{s}$) on the estimation errors and let $m$ grow to infinity at different rates relative to $p$. We consider four different growth rates with $m=3p^{1/2},3p^{2/3},3p^{3/4},1.8p$ and set $p=50,100,200,500$, $\delta=0,0.25$ and $n=1000$. In ${\mathbf{A}}^{s}$, each column has $m/3$ nonzero elements. Specifically, the first $m/3$ elements in its first column, the middle $m/3$ elements in the second column, and the last $m/3$ elements in the third column are nonzero, while all other elements are zero. Table 3 and Table 4 show the average and standard deviation of the estimation errors for the loading space and the sparsity level, respectively. It can be seen that our proposed approach outperforms other approaches. When $p$ is fixed, as $m$ increases, the estimation error first decreases and then increases, which aligns with our conclusion in Theorem 2. When the loading matrix is sparse, the estimation error is dominated by the bias; when the loading matrix is dense, it is dominated by the variance.
@@
$\delta$ $m$ $p$ eigen varimax1 varimax2 SO sparse
0 $3p^{1/2}$ 50 0.045(0.012) 0.038(0.013) 0.028(0.013) 0.383(0.173) 0.019(0.010)
$3p^{2/3}$ 50 0.029(0.007) 0.022(0.007) 0.040(0.011) 0.083(0.067) 0.016(0.006)
$3p^{3/4}$ 50 0.025(0.006) 0.024(0.005) 0.059(0.012) 0.055(0.028) 0.016(0.005)
$1.8p$ 50 0.020(0.004) 0.024(0.004) 0.086(0.011) 0.034(0.014) 0.017(0.004)
$3p^{1/2}$ 100 0.051(0.014) 0.037(0.015) 0.036(0.015) 0.477(0.142) 0.018(0.013)
$3p^{2/3}$ 100 0.033(0.007) 0.021(0.007) 0.060(0.011) 0.096(0.098) 0.016(0.004)
$3p^{3/4}$ 100 0.027(0.006) 0.016(0.004) 0.080(0.010) 0.047(0.023) 0.016(0.004)
$1.8p$ 100 0.019(0.004) 0.026(0.004) 0.134(0.010) 0.027(0.009) 0.016(0.004)
$3p^{1/2}$ 200 0.061(0.015) 0.037(0.016) 0.049(0.017) 0.563(0.034) 0.018(0.013)
$3p^{2/3}$ 200 0.037(0.007) 0.019(0.006) 0.083(0.010) 0.101(0.118) 0.015(0.003)
$3p^{3/4}$ 200 0.030(0.006) 0.016(0.004) 0.117(0.010) 0.043(0.023) 0.016(0.004)
$1.8p$ 200 0.019(0.004) 0.029(0.004) 0.213(0.009) 0.023(0.007) 0.016(0.003)
$3p^{1/2}$ 500 0.074(0.016) 0.034(0.011) 0.072(0.020) 0.576(0.004) 0.018(0.012)
$3p^{2/3}$ 500 0.043(0.008) 0.018(0.006) 0.133(0.010) 0.234(0.205) 0.016(0.003)
$3p^{3/4}$ 500 0.033(0.006) 0.018(0.004) 0.193(0.009) 0.039(0.021) 0.016(0.003)
$1.8p$ 500 0.019(0.004) 0.044(0.003) 0.392(0.009) 0.022(0.005) 0.016(0.003)
0.25 $3p^{1/2}$ 50 0.058(0.016) 0.052(0.018) 0.032(0.013) 0.518(0.092) 0.029(0.017)
$3p^{2/3}$ 50 0.041(0.010) 0.034(0.011) 0.043(0.011) 0.289(0.166) 0.025(0.009)
$3p^{3/4}$ 50 0.036(0.008) 0.034(0.008) 0.061(0.012) 0.181(0.117) 0.025(0.008)
$1.8p$ 50 0.030(0.007) 0.032(0.006) 0.089(0.011) 0.101(0.061) 0.027(0.006)
$3p^{1/2}$ 100 0.068(0.020) 0.057(0.023) 0.041(0.018) 0.564(0.025) 0.028(0.023)
$3p^{2/3}$ 100 0.049(0.010) 0.035(0.012) 0.062(0.011) 0.459(0.133) 0.025(0.007)
$3p^{3/4}$ 100 0.042(0.009) 0.028(0.010) 0.082(0.010) 0.281(0.165) 0.027(0.008)
$1.8p$ 100 0.032(0.007) 0.034(0.006) 0.136(0.010) 0.092(0.045) 0.029(0.007)
$3p^{1/2}$ 200 0.086(0.023) 0.065(0.026) 0.053(0.017) 0.575(0.005) 0.032(0.027)
$3p^{2/3}$ 200 0.057(0.011) 0.034(0.011) 0.086(0.010) 0.558(0.049) 0.026(0.007)
$3p^{3/4}$ 200 0.049(0.010) 0.028(0.009) 0.118(0.010) 0.465(0.130) 0.029(0.008)
$1.8p$ 200 0.035(0.007) 0.038(0.006) 0.215(0.009) 0.094(0.049) 0.031(0.006)
$3p^{1/2}$ 500 0.114(0.030) 0.071(0.035) 0.076(0.020) 0.577(0.002) 0.042(0.044)
$3p^{2/3}$ 500 0.072(0.014) 0.033(0.011) 0.135(0.010) 0.576(0.004) 0.031(0.011)
$3p^{3/4}$ 500 0.058(0.011) 0.030(0.007) 0.195(0.009) 0.572(0.008) 0.032(0.008)
$1.8p$ 500 0.039(0.008) 0.052(0.005) 0.392(0.009) 0.133(0.087) 0.035(0.007)

### v3

@@ 5 Simulation Study
In this section, we use four examples to illustrate the performance of our proposed method. The first three examples are used to compare our proposed approach with the method proposed by Lam et al. (2011), varimax rotation, and the method proposed by Wu et al. (2026). Since these methods adopt different ways to define the loading matrix, we cannot directly evaluate the estimation error of the loading matrix. Here we report the estimation error of the loading space for a fair comparison, $)\mbox{$\mathcal{D}$}({{\cal M}}({\mathbf{Q}}),{{\cal M}}(\widehat{{\mathbf{Q}}}$)), which is defined in (23). In all examples columns of $\mathbf{Q}$} are non-orthogonal. In Section 5.1, we set different values for $\delta$, $p$, $n$, and $m$ grows to infinity as fast as $p$. In Section 5.2, we allow $m$ to grow at different rates of $p$ to demonstrate the impact of $m$ on the estimation results. In Section 5.3, the sparsity level in each column of $\mathbf{Q}$} is different and we present the estimation of $\mathbf{Q}$} with our method. “eigen” represents the method by Lam et al. (2011), “SFM” represents the method in Wu et al. (2026) when the sparsity level is fixed at the true, and “sparse” represents our proposed method with tuning parameter selected based on BIC. For varimax rotation, to obtain sparse estimates, we set loadings with absolute values below a threshold to zero, and we consider two thresholds: 0.01 (“varmax1”) and 0.05 (“varimax2”).
Datasets are simulated from model (2). The nonzero elements in ${\mathbf{A}}^{s}$ are simulated from a truncated standard normal distribution with absolute values bounded above by 0.1 divided by $(m/r)^{\delta/2}$. We set $r=3$, and $\mathbf{f}_{t}^{s}$ is generated from three independent AR(1) processes with AR coefficient of $0.9$ and innovation variance of 1. The diagonal elements of $\boldsymbol{\Sigma}_{e,t}$ are all 1, and its off-diagonal elements are 0.5. The number of factors is assumed to be known. For each setting, we generate 300 samples.
@@ 5.1 Study on $\delta$, $p$ and $n$
In this example, we consider different combinations of $p$ and $n$, where $p=20,50,100,200,500$ and $n=100,200,500$. In ${\mathbf{A}}^{s}$, the first $0.4p$ elements in the first column, and the last $0.4p$ elements in the third column are nonzero, while all other elements in these two columns are zero. In the second column, the first $0.3p$ elements are zero, the following $0.4p$ elements are nonzero, and the remaining $0.3p$ elements are zero. Note that the second column of ${\mathbf{A}}^{s}$ contains nonzero elements that overlap with those in both the first and third columns, making the whole matrix non-orthogonal. Figure 1 gives an example of the simulated loading matrix when $p=20$. To illustrate the difference between our settings and the ones in Wu et al. (2026), we apply singular value decomposition to $\mathbf{M}$} to obtain ${\mathbf{A}}^{\text{svd}}$, the loading matrix assumed in Wu et al. (2026), which are orthogonal. We also standardize the columns of ${\mathbf{A}}^{s}$ so that each column has unit $L_{2}$ norm. Figure 1 shows that ${\mathbf{A}}^{s}$, the loading matrix in our settings, is very sparse, while ${\mathbf{A}}^{\text{svd}}$ is not at all. Although many elements in ${\mathbf{A}}^{\text{svd}}$ have small magnitude, thy are not exactly zeros.
Figure 1: An example of structures of loading matrices when $p=20$.
Table 1 shows the average estimation error of the loading space for different methods with $\delta=0$ and $\delta=0.25$. Our proposed estimator outperforms others, especially when the sample size is small. It is interesting that when the threshold for the varimax is appropriately selected, varimax performs better than the method of Lam et al. (2011), which does not account for the sparsity of the loading matrix. However, the optimal threshold value depends on $p$, $n$, and $\delta$, so selecting the threshold value is particularly challenging when analyzing real data.
@@
$\delta$ $n$ $p$ eigen varimax1 varimax2 SFM sparse
0 100 20 0.115(0.056) 0.115(0.056) 0.107(0.057) 0.346(0.213) 0.095(0.057)
100 50 0.110(0.038) 0.109(0.038) 0.103(0.033) 0.242(0.180) 0.088(0.036)
100 100 0.106(0.031) 0.103(0.032) 0.120(0.019) 0.182(0.125) 0.084(0.028)
100 200 0.103(0.028) 0.093(0.030) 0.174(0.012) 0.152(0.084) 0.082(0.024)
100 500 0.104(0.028) 0.085(0.028) 0.310(0.011) 0.144(0.071) 0.084(0.024)
200 20 0.066(0.022) 0.066(0.023) 0.066(0.021) 0.274(0.224) 0.052(0.023)
200 50 0.062(0.016) 0.060(0.017) 0.076(0.013) 0.127(0.114) 0.047(0.014)
200 100 0.061(0.015) 0.056(0.016) 0.105(0.010) 0.121(0.077) 0.046(0.013)
200 200 0.061(0.015) 0.050(0.015) 0.164(0.010) 0.118(0.064) 0.046(0.012)
200 500 0.061(0.014) 0.049(0.010) 0.303(0.009) 0.114(0.064) 0.046(0.012)
500 20 0.035(0.010) 0.035(0.010) 0.047(0.013) 0.208(0.215) 0.026(0.010)
500 50 0.035(0.008) 0.033(0.008) 0.066(0.012) 0.091(0.091) 0.024(0.007)
500 100 0.035(0.008) 0.032(0.008) 0.099(0.010) 0.081(0.056) 0.024(0.006)
500 200 0.035(0.007) 0.029(0.006) 0.159(0.009) 0.086(0.061) 0.024(0.006)
500 500 0.035(0.007) 0.035(0.004) 0.301(0.009) 0.090(0.063) 0.024(0.006)
0.25 100 20 0.152(0.074) 0.152(0.074) 0.143(0.076) 0.362(0.203) 0.130(0.077)
100 50 0.164(0.063) 0.163(0.063) 0.147(0.065) 0.276(0.172) 0.139(0.065)
100 100 0.170(0.053) 0.168(0.054) 0.159(0.046) 0.232(0.126) 0.145(0.053)
100 200 0.178(0.051) 0.173(0.053) 0.204(0.033) 0.211(0.094) 0.154(0.049)
100 500 0.203(0.058) 0.191(0.062) 0.334(0.023) 0.225(0.086) 0.181(0.055)
200 20 0.087(0.030) 0.087(0.030) 0.082(0.030) 0.282(0.218) 0.072(0.031)
200 50 0.091(0.023) 0.089(0.024) 0.089(0.018) 0.149(0.115) 0.073(0.022)
200 100 0.097(0.024) 0.093(0.025) 0.117(0.014) 0.147(0.080) 0.078(0.021)
200 200 0.106(0.025) 0.096(0.027) 0.174(0.012) 0.147(0.056) 0.087(0.021)
200 500 0.118(0.028) 0.098(0.029) 0.312(0.011) 0.153(0.055) 0.100(0.023)
500 20 0.046(0.013) 0.046(0.013) 0.053(0.014) 0.211(0.212) 0.036(0.014)
500 50 0.051(0.012) 0.049(0.012) 0.071(0.012) 0.100(0.089) 0.038(0.011)
500 100 0.056(0.013) 0.050(0.013) 0.104(0.011) 0.095(0.055) 0.042(0.011)
500 200 0.060(0.013) 0.049(0.013) 0.163(0.009) 0.102(0.055) 0.046(0.011)
500 500 0.067(0.014) 0.053(0.010) 0.304(0.009) 0.110(0.056) 0.052(0.011)
Table 2 shows the average value of |mm^||m-$\widehat${m}|, with the standard deviation given in parentheses, where $\widehat{m}$ is the number of nonzero elements in the estimated loading matrix. Overall, our proposed method effectively recovers the sparsity of the loading matrix. The performance of “sparse” is much better than other methods. Although varimax rotation may be able to estimate the loading space accurately when the threshold is appropriately selected, it cannot recover the sparsity as effectively as our proposed method.
@@
Table 2: Mean and standard deviation (in parentheses) of |mm^||m-$\widehat${m}| for the example in Section 5.1
@@
$n$ $p$ varimax1 varimax2 SFM sparse varimax1 varimax2 SFM sparse
100 20 26.1(3.5) 4.4(3.2) 27.9(7.0) 2.6(3.1) 27.7(3.2) 6.1(3.9) 27.7(6.9) 3.4(3.8)
100 50 49.6(9.1) 5.3(2.9) 70.5(17.5) 5.0(6.9) 58.8(8.1) 5.2(4.9) 69.6(17.6) 7.0(8.3)
100 100 65.8(20.1) 22.2(4.1) 143.3(37.0) 7.8(12.3) 96.0(18.9) 20.5(5.3) 141.7(37.3) 12.9(19.3)
100 200 63.3(34.9) 72.8(6.1) 282.8(73.0) 10.7(14.8) 140.2(39.4) 72.4(7.4) 279.1(72.4) 21.5(28.6)
100 500 31.0(40.7) 300.1(9.3) 731.1(183.6) 19.3(29.9) 195.3(103.0) 302.6(9.3) 718.6(184.9) 64.2(70.9)
200 20 22.2(4.5) 2.6(2.5) 29.1(7.1) 2.3(3.5) 24.2(4.1) 3.2(2.8) 29.1(7.0) 3.0(3.9)
200 50 33.9(10.6) 6.2(2.6) 75.7(15.5) 2.2(3.3) 44.5(9.9) 5.8(2.6) 75.4(15.3) 4.0(6.0)
200 100 34.9(17.3) 22.6(3.8) 144.8(35.2) 3.2(5.4) 62.1(19.7) 22.6(3.9) 143.9(35.2) 5.9(10.0)
200 200 18.0(16.8) 72.4(6.4) 284.7(73.0) 5.6(9.2) 68.2(34.9) 72.6(6.3) 282.8(72.6) 11.4(20.2)
200 500 23.9(6.0) 298.6(9.7) 723.1(185.4) 9.6(13.8) 41.5(49.8) 299.6(9.7) 716.8(184.5) 26.8(31.0)
500 20 18.2(5.3) 2.0(2.0) 30.5(6.3) 1.4(2.9) 20.2(4.9) 2.2(2.1) 30.4(6.3) 2.1(3.7)
500 50 22.0(9.4) 6.1(2.5) 77.4(15.1) 1.2(3.6) 29.5(9.5) 6.0(2.6) 77.3(15.1) 2.5(5.3)
500 100 17.1(12.4) 23.0(4.0) 152.3(33.7) 1.4(3.8) 29.5(14.8) 23.0(4.1) 152.0(33.7) 3.3(7.9)
500 200 6.1(8.3) 72.1(6.3) 301.8(67.9) 2.1(5.2) 16.6(15.7) 72.3(6.5) 300.7(67.9) 5.1(13.4)
500 500 24.6(4.5) 299.0(9.4) 745.5(175.8) 5.2(13.0) 23.6(6.1) 299.5(9.5) 743.1(174.7) 7.9(18.4)
@@ 5.2 Study on $m$
In this section, we evaluate the impact of $m$ (the total number of nonzero loadings in ${\mathbf{A}}^{s}$) on the estimation errors and let $m$ grow to infinity at different rates relative to $p$. We consider four different growth rates with $m=3p^{1/2},3p^{2/3},3p^{3/4},1.8p$ and set $p=50,100,200,500$, $\delta=0,0.25$ and $n=1000$. In ${\mathbf{A}}^{s}$, each column has $m/3$ nonzero elements. Specifically, the first $m/3$ elements in its first column, the last $m/3$ elements in the third column are nonzero, while all other elements are zero. In the second column, the first $p/3$ elements are zero, the following $m/3$ elements are nonzero, and the left elements are also zero. Table 3 and Table 4 show the average and standard deviation of the estimation errors for the loading space and the sparsity level, respectively. When $p$ is fixed, as $m$ increases, the estimation error first decreases and then increases, which aligns with our conclusion in Theorem 2. When the loading matrix is sparse, the estimation error is dominated by the bias; when the loading matrix is dense, it is dominated by the variance.
@@
$m$ $p$ $\delta=0$ $\delta=0.25$
$3p^{1/2}$ 50 0.019(0.010) 0.029(0.017)
$3p^{2/3}$ 50 0.016(0.006) 0.025(0.009)
$3p^{3/4}$ 50 0.016(0.005) 0.025(0.008)
$1.8p$ 50 0.017(0.004) 0.027(0.006)
$3p^{1/2}$ 100 0.018(0.013) 0.028(0.023)
$3p^{2/3}$ 100 0.016(0.004) 0.025(0.007)
$3p^{3/4}$ 100 0.016(0.004) 0.027(0.008)
$1.8p$ 100 0.016(0.004) 0.029(0.007)
$3p^{1/2}$ 200 0.018(0.013) 0.032(0.027)
$3p^{2/3}$ 200 0.015(0.003) 0.026(0.007)
$3p^{3/4}$ 200 0.016(0.004) 0.029(0.008)
$1.8p$ 200 0.016(0.003) 0.031(0.006)
$3p^{1/2}$ 500 0.018(0.012) 0.042(0.044)
$3p^{2/3}$ 500 0.016(0.003) 0.031(0.011)
$3p^{3/4}$ 500 0.016(0.003) 0.032(0.008)
$1.8p$ 500 0.016(0.003) 0.035(0.007)